# The Covariant Stone-von Neumann Theorem for Locally Compact Quantum Groups

### v1

@@ 2.4. Coaction Crossed Products
The second condition is variously referred to as continuity, coaction nondegeneracy, or the Podleś condition in the literature. The data $(\mathbb{G},A,\alpha)$ is also called a left coaction, or otherwise a left dynamical system. One analogously defines a right coaction of $\mathbb{G}$ on $A$ by swapping the positions of $A$ and $\mathbb{G}$ in the codomain of $\alpha$; as above, left coactions of $\mathbb{G}$ on $A$ correspond to right coactions of $\mathbb{G}^{\textnormal{op}}$ on $A$. Notice that the comultiplication of a locally compact quantum group is an example of a left (and right) coaction of $\mathbb{G}$ on itself. Another example is given by the (left) trivial coaction of $\mathbb{G}$ on $A$, denoted $(\mathbb{G},A,\text{triv})$. Here, $\text{triv}(a)=1\otimes a$.
Remark.
In the classical setting of locally compact groups, right coactions appear as the standard choice. In the quantum setting, however, left corepresentations and consequently left coactions appear to be more natural. Of course, in view of the comments above, the two notions are really more or less equivalent — the burden of working with right coactions is purely notational. Nevertheless, in this article, we will have use for both left and right coactions.
Given a left coaction $(\mathbb{G},A,\alpha)$ and any Hilbert $B$-module $X=X_{B}$, a covariant representation of $(\mathbb{G},A,\alpha)$ on $X_{B}$ is a pair of nondegenerate homomorphisms $(\mu,\pi)$, with $\mu\colon C_{0}^{u}(\widehat{\mathbb{G}})\to\mathcal{L}(X_{B})$ and $\pi\colon A\to\mathcal{L}(X_{B})$, subject to the condition that
@@
for all $a\in A$. Notice that the right hand side is conjugation by the left unitary corepresentation $(\text{id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})$. Elsewhere in the literature, a corepresentation is regarded as a pair comprising a homomorphism of the coefficient algebra and a left corepresentation on the common Hilbert module $X$. Given a right coaction $(A,\mathbb{G},\alpha)$, a covariant representation on $X=X_{B}$ is a pair $(\pi,m)$ with $\pi\colon A\to\mathcal{L}(X_{B})$ and $m\colon C_{0}^{u}(\widehat{\mathbb{G}}^{\textnormal{c}})\to\mathcal{L}(X_{B})$ subject to
@@
An important example of a covariant representation for $(\mathbb{G},A,\alpha)$ is given on the Hilbert module $L^{2}(\mathbb{G})\otimes A$ by the pair $(\hat{\lambda}\otimes 1,\alpha)$ – it is called the the regular representation of the coaction.
For the left coaction $(\mathbb{G},A,\alpha)$, the full crossed product is a $C^{*}$-algebra $\mathbb{G}\ltimes A$ together with a covariant representation
@@
For example, the crossed product for the trivial coaction $(\mathbb{G},A,\text{triv})$ is the maximal tensor product $C_{0}^{u}(\widehat{\mathbb{G}})\otimes_{\textnormal{max}}A$. For other fundamental examples, one constructs the full crossed product $\mathbb{G}\ltimes C_{0}(\mathbb{G})$ for the coaction $(\mathbb{G},C_{0}(\mathbb{G}),\Delta_{\mathbb{G}})$, and the reduced crossed product $\mathbb{G}\ltimes_{r}C_{0}(\mathbb{G})=(\hat{\lambda}\otimes 1\ltimes\Delta_{\mathbb{G}})(\mathbb{G}\ltimes C_{0}(\mathbb{G}))$. A locally compact quantum group $\mathbb{G}$ is regular if and only if $\mathbb{G}\ltimes_{r}C_{0}(\mathbb{G})\cong\mathbb{K}(L^{2}(\mathbb{G}))$. Similarly, $\mathbb{G}$ is called strongly regular provided that $\mathbb{G}\ltimes C_{0}(\mathbb{G})\cong\mathbb{K}(L^{2}(\mathbb{G}))$ ([venchilada, Definition 2.11]).
Let $(\mathbb{G},A,\alpha)$ be again an arbitrary left coaction. There is a left coaction $\widehat{\alpha}^{\rm op}$ of $\widehat{\mathbb{G}}^{\rm op}$ on $\mathbb{G}\ltimes A$, called the dual coaction to $\alpha$, uniquely characterized by
@@
Indeed, one may check that $((\hat{\lambda}\otimes j_{\mathbb{G}})\Delta_{\widehat{\mathbb{G}}^{\rm op}}^{u},1\otimes j_{A})$ forms a covariant representation of $(\mathbb{G},A,\alpha)$ on $C_{0}(\widehat{\mathbb{G}}^{\rm op})\otimes(\mathbb{G}\ltimes A)$. Then, by the universal property of $\mathbb{G}\ltimes A$, we obtain a non-degenerate ***-homomorphism $\widehat{\alpha}^{\rm op}=(\hat{\lambda}\otimes j_{\mathbb{G}})\Delta^{u}_{\widehat{\mathbb{G}}^{\rm op}}\ltimes(1\otimes j_{A})\colon\mathbb{G}\ltimes A\rightarrow M(C_{0}(\widehat{\mathbb{G}}^{\rm op})\otimes(\mathbb{G}\ltimes A))$ acting via (2).
@@
Let us end this section with a definition: for a coaction $(\mathbb{G},A,\alpha)$ with $\mathbb{G}$ regular, there is a canonical surjection $\widehat{\mathbb{G}}^{\rm op}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{K}(L^{2}(\mathbb{G}))\otimes A,$ which is given by the integrated form of the covariant representation $(\lambda^{c}\otimes 1,(\hat{\lambda}\otimes 1)\ltimes\alpha)$; the coaction is called maximal provided this surjection is an isomorphism.

### v2

@@ 2.4. Crossed Products
The second condition is variously referred to as continuity, coaction nondegeneracy, or the Podleś condition in the literature. The data $(\mathbb{G},A,\alpha)$ is also called a left dynamical system. One analogously defines a right action of $\mathbb{G}$ on $A$ by swapping the positions of $A$ and $\mathbb{G}$ in the codomain of $\alpha$; as above, left actions of $\mathbb{G}$ on $A$ correspond to right actions of $\mathbb{G}^{\textnormal{op}}$ on $A$. The triple $(A,\mathbb{G},\alpha)$ denotes a right dynamical system. Notice that the comultiplication of a locally compact quantum group is an example of a left (and right) action of $\mathbb{G}$ on itself. Another example is given by the (left) trivial action of $\mathbb{G}$ on $A$, denoted $(\mathbb{G},A,\text{triv})$. Here, $\text{triv}(a)=1\otimes a$.
Given a left dynamical system $(\mathbb{G},A,\alpha)$ and any Hilbert $B$-module $X=X_{B}$, a covariant representation of $(\mathbb{G},A,\alpha)$ on $X_{B}$ is a pair of nondegenerate homomorphisms $(\mu,\pi)$, with $\mu\colon C_{0}^{u}(\widehat{\mathbb{G}})\to\mathcal{L}(X_{B})$ and $\pi\colon A\to\mathcal{L}(X_{B})$, subject to the condition that
@@
for all $a\in A$. Notice that the right hand side is conjugation by the left unitary representation $(\text{id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})$. Elsewhere in the literature, a covariant representation is regarded as a pair comprising a homomorphism of the coefficient algebra and a left representation on the common Hilbert module $X$. Given a right dynamical system $(A,\mathbb{G},\alpha)$, a covariant representation on $X=X_{B}$ is a pair $(\pi,m)$ with $\pi\colon A\to\mathcal{L}(X_{B})$ and $m\colon C_{0}^{u}(\widehat{\mathbb{G}}^{\textnormal{c}})\to\mathcal{L}(X_{B})$ subject to
@@
An important example of a covariant representation for $(\mathbb{G},A,\alpha)$ is given on the Hilbert module $L^{2}(\mathbb{G})\otimes A$ by the pair $(\hat{\lambda}\otimes 1,\alpha)$ – it is called the the regular representation of the dynamical system.
For the left dynamical system $(\mathbb{G},A,\alpha)$, the full crossed product is a $C^{*}$-algebra $\mathbb{G}\ltimes A$ together with a covariant representation
@@
For example, the crossed product for the trivial action $(\mathbb{G},A,\text{triv})$ is the maximal tensor product $C_{0}^{u}(\widehat{\mathbb{G}})\otimes_{\textnormal{max}}A$. For other fundamental examples, one constructs the full crossed product $\mathbb{G}\ltimes C_{0}(\mathbb{G})$ for the dynamical system $(\mathbb{G},C_{0}(\mathbb{G}),\Delta_{\mathbb{G}})$, and the reduced crossed product $\mathbb{G}\ltimes_{r}C_{0}(\mathbb{G})=(\hat{\lambda}\otimes 1\ltimes\Delta_{\mathbb{G}})(\mathbb{G}\ltimes C_{0}(\mathbb{G}))$. A locally compact quantum group $\mathbb{G}$ is regular if and only if $\mathbb{G}\ltimes_{r}C_{0}(\mathbb{G})\cong\mathbb{K}(L^{2}(\mathbb{G}))$. Similarly, $\mathbb{G}$ is called strongly regular provided that $\mathbb{G}\ltimes C_{0}(\mathbb{G})\cong\mathbb{K}(L^{2}(\mathbb{G}))$ ([venchilada, Definition 2.11]). We remark that we obtain the same property, if we assume that these isomorphisms are implemented by canonical maps, or just abstract.
Let $(\mathbb{G},A,\alpha)$ be again an arbitrary left dynamical system. There is a left action $\widehat{\alpha}^{\rm op}$ of $\widehat{\mathbb{G}}^{\rm op}$ on $\mathbb{G}\ltimes A$, called the dual action to $\alpha$, uniquely characterized by
@@
Indeed, one may check that $((\hat{\lambda}\otimes j_{\mathbb{G}})\Delta_{\widehat{\mathbb{G}}^{\rm op}}^{u},1\otimes j_{A})$ forms a covariant representation of $(\mathbb{G},A,\alpha)$ on $C_{0}(\widehat{\mathbb{G}}^{\rm op})\otimes(\mathbb{G}\ltimes A)$. Then, by the universal property of $\mathbb{G}\ltimes A$, we obtain a non-degenerate *-homomorphism $\widehat{\alpha}^{\rm op}=(\hat{\lambda}\otimes j_{\mathbb{G}})\Delta^{u}_{\widehat{\mathbb{G}}^{\rm op}}\ltimes(1\otimes j_{A})\colon\mathbb{G}\ltimes A\rightarrow M(C_{0}(\widehat{\mathbb{G}}^{\rm op})\otimes(\mathbb{G}\ltimes A))$ acting via (2).
@@
For a dynamical system $(\mathbb{G},A,\alpha)$ with $\mathbb{G}$ regular, there is a canonical surjection $\widehat{\mathbb{G}}^{\rm op}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{K}(L^{2}(\mathbb{G}))\otimes A,$ which is given by the integrated form of the covariant representation $(\lambda^{c}\otimes 1,(\hat{\lambda}\otimes 1)\ltimes\alpha)$; the action $\alpha$ is called maximal provided this surjection is an isomorphism [venchilada, Definition 2.13].
We are now prepared to present a result of general interest to the field of locally compact quantum groups.
Proposition 2.2.
Let $\mathbb{G}$ be a locally compact quantum group. We have the following isomorphisms [^1] We are grateful to Stefaan Vaes for pointing us towards the proof of isomorphism $\mathbb{G}\ltimes C_{0}^{u}(\mathbb{G})\cong\mathbb{G}\ltimes C_{0}(\mathbb{G})$.:
  1. (1)
    $\mathbb{G}\ltimes C_{0}^{u}(\mathbb{G})\cong\widehat{\mathbb{G}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}})$,
  2. (2)
    $\mathbb{G}^{\textnormal{op}}\ltimes C_{0}^{u}(\mathbb{G}^{\textnormal{op}})\cong(\mathbb{G}\ltimes C_{0}^{u}(\mathbb{G}))^{\textnormal{op}},$
  3. (3)
    $\mathbb{G}\ltimes C_{0}^{u}(\mathbb{G})\cong\mathbb{G}\ltimes C_{0}(\mathbb{G})$.
Assume that $\mathbb{G}$ (equivalently: $\mathbb{G}^{\textnormal{op}}$ or $\widehat{\mathbb{G}}$) is regular. As a consequence we conclude that the following properties are equivalent:
  • $\mathbb{G}$ is strongly regular,
  • $\mathbb{G}^{\textnormal{op}}$ is strongly regular,
  • $\widehat{\mathbb{G}}$ is strongly regular,
  • trivial action of $\mathbb{G}$ on $\mathbb{C}$ is maximal.
Let us remark that each of the above isomorphisms is implemented by natural maps.
Proof.
Let $j_{\mathbb{G}}$ and $j_{C_{0}^{u}(\mathbb{G})}$ denote the canonical maps for the full crossed product $\mathbb{G}\ltimes C_{0}^{u}(\mathbb{G})$, and let $j_{\widehat{\mathbb{G}}}$ and $j_{C_{0}^{u}(\widehat{\mathbb{G}})}$ be the analogous maps for $\widehat{\mathbb{G}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}})$. We claim that $(j_{C_{0}^{u}(\widehat{\mathbb{G}})},j_{\widehat{\mathbb{G}}})$ is a covariant representation of the dynamical system $(\mathbb{G},C_{0}^{u}(\mathbb{G}),(\lambda\otimes\text{id})\Delta^{u})$ on $M(\widehat{\mathbb{G}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}}))$. Indeed, since $(j_{\widehat{\mathbb{G}}},j_{C_{0}^{u}(\widehat{\mathbb{G}})})$ forms a covariant representation of $(\widehat{\mathbb{G}},C_{0}^{u}(\widehat{\mathbb{G}}),(\widehat{\lambda}\otimes{\rm id})\widehat{\Delta}^{u})$, for $x^{u}=(\omega\otimes\text{id})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}})$, $\omega\in L^{1}(\widehat{\mathbb{G}})$, we have that
$\begin{split}&\quad\;(\text{id}\otimes j_{C_{0}^{u}(\widehat{\mathbb{G}})})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})^{*}(1\otimes j_{\widehat{\mathbb{G}}}(x^{u}))(\text{id}\otimes j_{C_{0}^{u}(\widehat{\mathbb{G}})})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})\\ &=(\omega\otimes\text{id}\otimes\text{id})\bigl{(}(\text{id}\otimes j_{C_{0}^{u}(\widehat{\mathbb{G}})})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})^{*}_{23}(\text{id}\otimes j_{\widehat{\mathbb{G}}})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}})_{13}(\text{id}\otimes j_{C_{0}^{u}(\widehat{\mathbb{G}})})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{23}\bigr{)}\\ &=(\text{id}\otimes\omega\otimes\text{id})\bigl{(}(\text{id}\otimes j_{\widehat{\mathbb{G}}})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}})_{23}(\text{id}\otimes j_{\widehat{\mathbb{G}}})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}})_{23}^{*}(\text{id}\otimes j_{C_{0}^{u}(\widehat{\mathbb{G}})})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})^{*}_{13}(\text{id}\otimes j_{\widehat{\mathbb{G}}})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}})_{23}\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad(\text{id}\otimes j_{C_{0}^{u}(\widehat{\mathbb{G}})})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{13}\bigr{)}\\ &=(\text{id}\otimes\omega\otimes\text{id})\bigl{(}(\text{id}\otimes j_{\widehat{\mathbb{G}}})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}})_{23}(\text{id}\otimes\text{id}\otimes j_{C_{0}^{u}(\widehat{\mathbb{G}})})(\text{id}\otimes(\widehat{\lambda}\otimes\text{id})\widehat{\Delta}^{u})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}*})(\text{id}\otimes j_{C_{0}^{u}(\widehat{\mathbb{G}})})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{13}\bigr{)}\\ &=(\text{id}\otimes\omega\otimes\text{id})\bigl{(}(\text{id}\otimes j_{\widehat{\mathbb{G}}})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}})_{23}\mathrm{W}^{\mathbb{G}*}_{12}(\text{id}\otimes j_{C_{0}^{u}(\widehat{\mathbb{G}})})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}*})_{13}(\text{id}\otimes j_{C_{0}^{u}(\widehat{\mathbb{G}})})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{13}\bigr{)}\\ &=(\text{id}\otimes\omega\otimes\text{id})\bigl{(}(\text{id}\otimes j_{\widehat{\mathbb{G}}})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}})_{23}\mathrm{W}^{\mathbb{G}*}_{12}\bigr{)}=(\omega\otimes\text{id}\otimes\text{id})\bigl{(}(\text{id}\otimes j_{\widehat{\mathbb{G}}})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}})_{13}\mathrm{W}^{\widehat{\mathbb{G}}}_{12}\bigr{)}\\ &=(\omega\otimes\text{id}\otimes\text{id})\bigl{(}(\text{id}\otimes\text{id}\otimes j_{\widehat{\mathbb{G}}})(\text{id}\otimes(\lambda\otimes\text{id})\Delta^{u})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}})\bigr{)}=(\text{id}\otimes j_{\widehat{\mathbb{G}}})((\lambda\otimes\text{id})\Delta^{u}(x^{u})),\end{split}$
which, by continuity of the maps involved, proves the claim. Hence, there is a unique non-degenerate *-homomorphism $j_{C_{0}^{u}(\widehat{\mathbb{G}})}\ltimes j_{\widehat{\mathbb{G}}}$: $\mathbb{G}\ltimes C_{0}^{u}(\mathbb{G})\rightarrow M(\widehat{\mathbb{G}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}}))$ such that
(3) $(j_{C_{0}^{u}(\widehat{\mathbb{G}})}\ltimes j_{\widehat{\mathbb{G}}})j_{\mathbb{G}}=j_{C_{0}^{u}(\widehat{\mathbb{G}})},\quad(j_{C_{0}^{u}(\widehat{\mathbb{G}})}\ltimes j_{\widehat{\mathbb{G}}})j_{C_{0}^{u}(\mathbb{G})}=j_{\widehat{\mathbb{G}}}.$
By duality, we obtain a non-degenerate *-homomorphism $j_{C_{0}^{u}(\mathbb{G})}\ltimes j_{\mathbb{G}}\colon\widehat{\mathbb{G}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}})\rightarrow M(\mathbb{G}\ltimes C_{0}^{u}(\mathbb{G}))$ such that
(4) $(j_{C_{0}^{u}(\mathbb{G})}\ltimes j_{\mathbb{G}})j_{\widehat{\mathbb{G}}}=j_{C_{0}^{u}(\mathbb{G})},\quad(j_{C_{0}^{u}(\mathbb{G})}\ltimes j_{\mathbb{G}})j_{C_{0}^{u}(\widehat{\mathbb{G}})}=j_{\mathbb{G}}.$
Putting together (3) and (4), we get
$(j_{C_{0}^{u}(\mathbb{G})}\ltimes j_{\mathbb{G}})(j_{C_{0}^{u}(\widehat{\mathbb{G}})}\ltimes j_{\widehat{\mathbb{G}}})j_{\mathbb{G}}=(j_{C_{0}^{u}(\mathbb{G})}\ltimes j_{\mathbb{G}})j_{C_{0}^{u}(\widehat{\mathbb{G}})}=j_{\mathbb{G}}$
and
$(j_{C_{0}^{u}(\mathbb{G})}\ltimes j_{\mathbb{G}})(j_{C_{0}^{u}(\widehat{\mathbb{G}})}\ltimes j_{\widehat{\mathbb{G}}})j_{C_{0}^{u}(\mathbb{G})}=(j_{C_{0}^{u}(\mathbb{G})}\ltimes j_{\mathbb{G}})j_{\widehat{\mathbb{G}}}=j_{C_{0}^{u}(\mathbb{G})},$
which shows that $j_{C_{0}^{u}(\widehat{\mathbb{G}})}\ltimes j_{\widehat{\mathbb{G}}}$ is an isomorphism with inverse $j_{C_{0}^{u}(\mathbb{G})}\ltimes j_{\mathbb{G}}$. This shows the first claimed isomorphim.
Isomorphism $(2)$ can be proved by a similar technique as above, combining the canonical maps into crossed product, universal property of full crossed product together with canonical antimultiplicative maps: unitary antipode, $C_{0}^{u}(\mathbb{H})\rightarrow C_{0}^{u}(\mathbb{H}^{c})$ and $A\rightarrow A^{\textnormal{op}}$ (for LCQGs $\mathbb{H}$ and $C^{*}$-algebras $A$).
To see isomorphism $(3)$, let us first make an auxiliary observation. Consider arbitrary representation (i.e. a non-degenerate *-homomorphism) $\mathbb{G}\ltimes C_{0}^{u}(\mathbb{G})\rightarrow\mathbb{B}(H)$; it is given by integration $\mu\ltimes\pi$ of a covariant representation $(\mu,\pi)$ of $(\mathbb{G},C_{0}^{u}(\mathbb{G}),(\lambda\otimes\text{id})\Delta^{u})$ on $H$. The covariance condition reads
(5) $(\text{id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})^{*}(1\otimes\pi(y^{u}))(\text{id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})=(\text{id}\otimes\pi)\circ(\lambda\otimes\text{id})\Delta^{u}(y^{u})$
for $y^{u}\in C_{0}^{u}(\mathbb{G})$. In particular, for $y^{u}=(\text{id}\otimes\widehat{\omega})\mathds{W}^{\mathbb{G}}$ with $\widehat{\omega}\in L^{1}(\widehat{\mathbb{G}})$ we obtain
$\begin{split}&\quad\;(\text{id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})^{*}(1\otimes\pi(y^{u}))(\text{id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})=(\text{id}\otimes\text{id}\otimes\widehat{\omega})\bigl{(}(\lambda\otimes\pi\otimes\text{id})(\Delta^{u}\otimes\text{id})(\mathds{W}^{\mathbb{G}})\bigr{)}\\ &=(\text{id}\otimes\text{id}\otimes\widehat{\omega})\bigl{(}(\text{id}\otimes\pi\otimes\text{id})({\text{\reflectbox{$\mathds{V}$}}\>\!}^{\mathbb{G}}_{12}\mathrm{W}^{\mathbb{G}}_{13}\text{\reflectbox{$\mathds{V}$}}\>\!^{\mathbb{G}*}_{12})\bigr{)}=(\text{id}\otimes\text{id}\otimes\widehat{\omega})\bigl{(}(\text{id}\otimes\pi)(\text{\reflectbox{$\mathds{V}$}}\>\!^{\mathbb{G}})_{12}\mathrm{W}^{\mathbb{G}}_{13}(\text{id}\otimes\pi)(\text{\reflectbox{$\mathds{V}$}}\>\!^{\mathbb{G}})^{*}_{12}\bigr{)}\\ &=(\text{id}\otimes\pi)(\text{\reflectbox{$\mathds{V}$}}\>\!^{\mathbb{G}})((\text{id}\otimes\widehat{\omega})(\mathrm{W}^{\mathbb{G}})\otimes 1)(\text{id}\otimes\pi)(\text{\reflectbox{$\mathds{V}$}}\>\!^{\mathbb{G}})^{*}=(\text{id}\otimes\pi)(\text{\reflectbox{$\mathds{V}$}}\>\!^{\mathbb{G}})(\lambda(y^{u})\otimes 1)(\text{id}\otimes\pi)(\text{\reflectbox{$\mathds{V}$}}\>\!^{\mathbb{G}})^{*}.\end{split}$
By continuity, the above equation holds for all $y^{u}\in C_{0}^{u}(\mathbb{G})$. Consequently, we see that $\pi$ vanishes on the kernel of $\lambda$ and there is a non-degenerate *-homomorphism $\widetilde{\pi}\colon C_{0}(\mathbb{G})\rightarrow\mathbb{B}(H)$ such that $\pi=\widetilde{\pi}\circ\lambda$. One immediately sees from (5) that $(\mu,\widetilde{\pi})$ forms a covariant representation of $(\mathbb{G},C_{0}(\mathbb{G}),\Delta)$, hence we obtain non-degenerate *-homomorphism $\mu\ltimes\widetilde{\pi}\colon\mathbb{G}\ltimes C_{0}(\mathbb{G})\rightarrow\mathbb{B}(H)$ such that $(\mu\ltimes\widetilde{\pi})(j_{\mathbb{G}}^{r}\ltimes j_{C_{0}(\mathbb{G})}\lambda)=\mu\ltimes\pi$, where $j_{\mathbb{G}}^{r},j_{C_{0}(\mathbb{G})}$ are the canonical maps for the crossed product $\mathbb{G}\ltimes C_{0}(\mathbb{G})$.
We can apply this reasoning to $(\mu,\pi)=(j_{\mathbb{G}},j_{C_{0}^{u}(\mathbb{G})})$. Then we obtain map $j_{\mathbb{G}}\ltimes\widetilde{j}_{C_{0}^{u}(\mathbb{G})}\colon\mathbb{G}\ltimes C_{0}(\mathbb{G})\rightarrow\mathbb{G}\ltimes C_{0}^{u}(\mathbb{G})$ satisfying
$(j_{\mathbb{G}}\ltimes\widetilde{j}_{C_{0}^{u}(\mathbb{G})})(j_{\mathbb{G}}^{r}\ltimes j_{C_{0}(\mathbb{G})}\lambda)=j_{\mathbb{G}}\ltimes j_{C_{0}^{u}(\mathbb{G})}=\text{id}.$
This prove that the (clearly surjective) non-degenerate *-homomorphism $j_{\mathbb{G}}^{r}\ltimes j_{C_{0}(\mathbb{G})}\lambda\colon\mathbb{G}\ltimes C_{0}^{u}(\mathbb{G})\rightarrow\mathbb{G}\ltimes C_{0}(\mathbb{G})$ is in fact an isomorphism.
Now, let $\mathbb{G}$ be a regular locally compact quantum group. Assume that the trivial action $\mathbb{C}\rightarrow M(C_{0}(\mathbb{G})\otimes\mathbb{C})$ is maximal, i.e. that the canonical surjection $\widehat{\mathbb{G}}^{\rm op}\ltimes(\mathbb{G}\ltimes\mathbb{C})\rightarrow\mathbb{K}(L^{2}(\mathbb{G}))$ is an isomorphism. Using isomorphisms proved above we have the following:
$\begin{split}\mathbb{K}(L^{2}(\mathbb{G}))&\cong\widehat{\mathbb{G}}^{\rm op}\ltimes(\mathbb{G}\ltimes\mathbb{C})=\widehat{\mathbb{G}}^{\rm op}\ltimes C_{0}^{u}(\widehat{\mathbb{G}})\cong\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}}^{\textnormal{op}})\cong(\mathbb{G}\ltimes C_{0}^{u}(\mathbb{G}))^{\textnormal{op}}\cong(\mathbb{G}\ltimes C_{0}(\mathbb{G}))^{\textnormal{op}}.\end{split}$
Since $\mathbb{K}(L^{2}(\mathbb{G}))^{\textnormal{op}}\cong\mathbb{K}(L^{2}(\mathbb{G}))$, it shows that $\mathbb{G}$ is strongly regular. This reasoning can easily be reversed, hence we see that if $\mathbb{G}$ is strongly regular, then $\mathbb{G}\curvearrowright\mathbb{C}$ is maximal. Equivalence of these properties with strong regularity of $\widehat{\mathbb{G}}$ follows from isomorphism $(1)$, and with strong regularity of $\mathbb{G}^{\textnormal{op}}$ from $(2)$. ∎
2.5. The Mackey–Stone–von Neumann Theorem for Quantum Groups
Various authors call the Stone-von Neumann theorem the isomorphism $G\ltimes C_{0}(\widehat{G})\cong\mathbb{K}(L^{2}(G))$ (see [raeburnwilliams, Theorem C.34] and [huneufangruan, Theorem 3.4] for a more general result in the quantum group setting) which is a crucial ingredient in Rieffel’s proof of Mackey’s version of the theorem. As far as the authors know, there is no Mackey–Stone–von Neumann theorem for quantum groups. In what follows, we formulate and prove a more general result relevant to Hilbert modules over any elementary $C^{*}$-algebra. Here, with the technology described above in hand, we discuss known related results. In particular, we highlight some parallels between our work, [RT18], and [Rieffel]. The authors of the former article consider quantum group-twisted tensor product $C^{*}$-algebras in the settings of maximal tensor products, building on the work in the minimal setting of [mrw]. Their gadget subsumes maximal tensor products, skew-commutative tensor products of $\mathbb{Z}/2\mathbb{Z}$-graded $C^{*}$-algebras, and the construction of importance to this discussion, universal crossed products. The details we need are found in [RT18]*Section 6.2.
Consider the dynamical system $(\widehat{\mathbb{G}}^{\textnormal{op}},C_{0}^{u}(\widehat{\mathbb{G}}),(\hat{\lambda}\otimes\text{id})\circ\Delta^{u}_{\widehat{\mathbb{G}}^{\textnormal{op}}})$ and a covariant representation $(\mu,\pi)$ for it on a Hilbert $\mathbb{C}$-module $X$ (a Hilbert space representation). In the proof of [RT18]*Theorem 6.2, the authors show (up to chirality of left-right representations) that this is equivalent to the pair furnishing a $\mathrm{W}^{\mathbb{G}}$-commutative representation, which entails commutation of the two representations modulo some interference with respect to $\mathrm{W}^{\mathbb{G}}$; c.f. our Definition 3.1 and Proposition 3.9. Now, if the trivial action of $\mathbb{G}$ on $\mathbb{C}$ is maximal (so that $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}})\cong\mathbb{K}(L^{2}(\mathbb{G}))$), then the integrated form $\mu\ltimes\pi$ implements a representation of the elementary $C^{*}$-algebra $\mathbb{K}(L^{2}(\mathbb{G}))$. Since all irreducible representations of $\mathbb{K}(L^{2}(\mathbb{G}))$ are unitarily equivalent to the identity representation, it follows that $\mu\ltimes\pi$ is a multiple of the identity representation. Upon disintegration of any irreducible subrepresentation, one may reinterpret the covariant subrepresentation as a $\mathrm{W}^{\mathbb{G}}$-commutative representation of it’s own. This is essentially the strategy Rieffel used in [Rieffel] to prove his version of the Mackey-Stone-von Neumann Theorem, and the principle behind our Theorem 4.3. Thus, maximality of the trivial action is a necessary condition. In Corollary 4.9, we show that it is also sufficient.